1. Assume a closed economy and no government. Also assume consumption C=50 + 0.8Yd and investment I=80

a) derive the equation for saving

b) find the equilibrium output

c) what is the implied multiplier


1
Expert's answer
2021-08-31T17:03:11-0400
"Solution"

"(a)\\
Y=C+I\\
C=50+0.8Y_d\\I=80\\
S=Y-C\\
S=(50+0.8Y_d+80)-50+0.8Y_d\\
S=130+0.8Y_d-50+0.8Y_d\\
S=80+1.6Y_d\\
(b)\\
EQ \ where\ Y=C+I\\
Y=50+0.8Y_d+80\\
Y=130+0.8Y_d\\
Y-0.8Y=130\\
0.2 Y=130\\
Y=650\\
(c)\\
m=\frac{1}{1-mpc}\\
m=\frac{1}{1-0.8}\\
m=\frac{1}{0.2}\\
m=5"